OpenStudy (anonymous):

2/3 of A is 5/12 and 3/4 of B is 8/9. Which one is larger, A or B?

OpenStudy (amistre64):

just solve for A and B

OpenStudy (amistre64):

2/3 A = 5/12 3/4 B = 8/9